OpenStudy (anonymous):

How can you find the zeros of x^5-3x^4-24x^3-72x^2-25x+75=0, since it cannot be factored?

OpenStudy (anonymous):

\[x^5-3x^4-24x^3-72x^2-25x+75=0\]

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Are you sure, it is not +72x² ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

could be a typo in the question but that is what it says

OpenStudy (solomonzelman):

well, for +72, \(\normalsize\color{black}{ x^5-3x^4-24x^3+72x^2-25x+75=0\huge\color{white}{ \rm │ }}\) \(\normalsize\color{black}{ x^4(x-3)-24x^2(x-3)-5(x-3)=0\huge\color{white}{ \rm │ }}\) \(\normalsize\color{black}{ (x^4-24x^2-5)(x-3)=0\huge\color{white}{ \rm │ }}\) and then you get `x-3` as the obvious part, and the first factor with completing the square....

OpenStudy (solomonzelman):

I don't know how to do it, with -72x²... sorry _:(

OpenStudy (anonymous):

OK, thanks anyways.

OpenStudy (anonymous):

@ganeshie8 @DangerousJesse

OpenStudy (solomonzelman):

I didn't help much -:(

OpenStudy (anonymous):

It's fine

OpenStudy (solomonzelman):

wolfram

OpenStudy (dangerousjesse):

Yep

OpenStudy (dangerousjesse):

I always use wolfram to graph

OpenStudy (anonymous):

Thats how I though you could do it but there are also 2 complex solutions

OpenStudy (anonymous):

*is

OpenStudy (solomonzelman):

I am really thinking it should be +72x²

OpenStudy (anonymous):

no, are

OpenStudy (solomonzelman):

Then it is much better and easier.

OpenStudy (solomonzelman):

yes there `are` also 2 complex solutions.

OpenStudy (dangerousjesse):

Complex solutions are fairly quick to find, have you already found the two?

OpenStudy (anonymous):

I'm not sure on how to find the complex solutions

OpenStudy (dangerousjesse):

\[x = \sqrt of x^5-3 x^4-24 x^3-72 x^2-25 x+75 \approx x = -1.84865-2.1615 i\]

OpenStudy (solomonzelman):

there are approximated just like on wolfram.

OpenStudy (solomonzelman):

You can use `~` for space.

OpenStudy (solomonzelman):

If you put ~~ it will give more space, and ~~~ even more space, and on...

OpenStudy (dangerousjesse):

These are the exact complex solutions

OpenStudy (dangerousjesse):

\[x = \sqrt of x^5-3 x^4-24 x^3-72 x^2-25 x+75 \approx x = -1.84865+2.1615 i\]

OpenStudy (anonymous):

OK, so use the graphing software to get the real solutions, and use that method to get the complex ones.

OpenStudy (dangerousjesse):

Yeah, I'd do that if I were you.

OpenStudy (anonymous):

OK, thanks, again :P

OpenStudy (dangerousjesse):

No problem.